# SlotMatch: Distilling Temporally Consistent Object-Centric Representations for Unsupervised Video Segmentation

### v2

@@ Introduction
MOVi-E.
The Multi-Object Video (MOVi-E) dataset (Ghorbani et al. 2021) is generated using the Kubric simulator, which provides ground-truth object segmentation for precise evaluation. MOVi-E includes up to 23 objects per scene and is filmed with linear camera motion. The dataset includes complex object interactions, occlusions, and realistic textures, and the videos have a 24-frame length and a $\times256$ resolution. We use the official train and validation splits, with 10,000 and 1,000 videos, respectively.
YTVIS-2021.
To evaluate scalability to real-world scenarios, we use the YouTube Video Instance Segmentation 2021 (YTVIS-2021) dataset (Yang et al. 2021b). It contains unconstrained real-world videos, capturing diverse scenes including indoor/outdoor environments, multiple object categories, and complex interactions. Videos vary in length (up to 76 frames). The official split has 2,985 training videos and 1,421 validation ones, which we resize to $\times518$.
@@
We conduct zero-shot experiments on Occluded Video Instance Segmentation (OVIS) (Qi et al. 2022), a dataset specifically focused on challenging scenarios. OVIS contains real-world videos with heavy occlusions, making it particularly suitable for testing temporal consistency when objects frequently disappear and reappear. The dataset features 607 training videos and 140 validation videos across 25 object categories.
Evaluation Metrics
We use two types of evaluation metrics, namely the Foreground Adjusted Rand Index (FG-ARI) and mean Best Overlap (mBO), and apply them at both image and video levels. FG-ARI measures how well the model groups pixels belonging to the same object. With a range of 0 to 1, and a higher value indicating better performance, its focus is on object discovery quality, i.e. if the model correctly identifies which pixels belong to the same object. FG-ARI compares predicted object masks against ground-truth segmentation masks. mBO measures the spatial precision of object masks, i.e. how accurately the predicted masks align with ground-truth boundaries. For each predicted mask, it finds the ground-truth mask with the highest IoU, then averages the IoU across all objects. mBO focuses on segmentation mask quality, measuring how precise and well-defined the predicted object boundaries are.
@@
Baselines
We compare SlotMatch against a range of recent object-centric video segmentation models. These include slot-based methods such as SAVi (Kipf et al. 2022), STEVE (Singh, Wu, and Ahn 2022), SOLV (Aydemir, Xie, and Guney 2023), VideoSAUR (Zadaianchuk, Seitzer, and Martius 2023), and its DINOv2-enhanced variant, VideoSAURv2 (Manasyan et al. 2025). We reproduce the results of SlotContrast (Manasyan et al. 2025), which serves as the teacher in our framework based on knowledge distillation. We further introduce several new baselines, each employing a different distillation strategy, as follows:
@@
  • SlotMatch (predicted) distills predicted slots (before slot attention) by minimizing cosine similarity.
Results
Quantitative results.
In Table 1, we present comparative results on MOVi-E and YTVIS-2021. Remarkably, SlotMatch consistently outperforms all prior methods across both datasets, including its teacher, SlotContrast (Manasyan et al. 2025). Despite having $3.6\times$ fewer parameters and nearly $2\times$ faster inference speed than SlotContrast (as per Table 2), our student achieves higher mask precision (mBO) and comparable or better object separation (FG-ARI). Notably, we improve mBO on YTVIS from 32.4 to 32.8, while reducing latency by nearly half, demonstrating that accurate object-centric segmentation does not require large models.
We find that mBO improvements are most consistent on the real-world YTVIS dataset, where the teacher may overfit to slot assignment, but underperform on precise masks. Interestingly, our student avoids this overfitting and better balances separation and alignment. This aligns with our hypothesis that slot supervision can offer better generalization than end-to-end training with larger capacity alone.
Qualitative results.
In Figure 3, we showcase qualitative comparisons of segmentation masks on challenging examples from MOVi-E and YTVIS. Our method produces sharper and more temporally-consistent masks than the student model without distillation. On MOVi-E, SlotMatch segments overlapping or partially occluded objects more cleanly, while on real-world YTVIS data, it shows improved boundary alignment and fewer slot collisions. These results highlight the benefit of slot-level supervision in guiding the student to focus on meaningful object structure.
Method #Params (M) $\downarrow$ #GFLOPs $\downarrow$ Inference Time (ms) $\uparrow$
Teacher 91.0 8825.7 347.4
SlotMatch 25.2 3266.5 186.0
@@ Results
Table 2: Comparison of inference speed, number of parameters, and GFLOPs between the SlotContrast teacher and our SlotMatch student. Measurements are performed on an NVIDIA A100 with 40GB VRAM, using video samples from the YTVIS-2021 dataset.
@@
Table 3: Zero-shot generalization results on the OVIS dataset, using models trained only on YTVIS. We compare the student trained from scratch (no KD) versus our SlotMatch student.
@@
Efficiency comparison.
[…]
Zero-shot generalization.
In Table 3, we compare SlotMatch against the student trained from scratch (without distillation), in the zero-shot setup on the challenging OVIS dataset. Notably, slot-match maintains competitive video-level performance (FG-ARI: 34.8 vs. 34.6) compared with the student trained from scratch, while preserving slightly better spatial mask quality (mBO: 25.5 vs. 24.9). This demonstrates that distilling structured slot representations enhances robustness to occlusion and domain shift.
Ablation Studies
We conduct ablations to isolate the contribution of each component in our framework. The ablation results are summarized in Tables 1 and 4.
Slot vs. feature and reconstruction KD.
To test whether slot representations are the best target for distillation, we compare our method with variants that distill encoder or decoder features via MSE (see Table 1). While feature KD and reconstruction KD bring modest gains over the student (no KD) baseline, they remain below our slot-based formulation, by significant margins. This supports our hypothesis that distilling structured object-centric representations leads to better temporal consistency and segmentation quality.
Cosine vs. MSE loss.
We compare the proposed loss based on cosine similarity with an alternative loss based on MSE over slot representations. Cosine-based distillation yields more stable improvements across datasets, especially on YTVIS, likely due to its scale invariance and better alignment with semantic structure in high-dimensional spaces.
Matching predicted vs. corrected slots.
Another ablation study targets the placement of our $\mathcal{L}_{\text{slot-KD}}$ loss, i.e. before or after slot attention. In general, it seems to be more beneficial to apply the distillation to corrected slots, after slot attention. Yet, temporal consistency (measured at the video level) can be accurately ensured by distilling predicted slots, as confirmed by the results on YTVIS.
With or without reconstruction KD.
@@
To empirically determine if the proposed $\mathcal{L}_{\text{slot-KD}}$ loss is sufficient or not, we carry out experiments with an enhanced version of SlotMatch, where the distillation is also applied over reconstructed features via the $\mathcal{L}_{\text{rec-KD}}$ loss. Perhaps surprisingly, this double distillation procedure degrades performance by considerable margins on MOVi-E. The results show that adding additional distillation losses can make the distillation more complex and difficult to tune, eventually leading to inferior results. The empirical results only confirm Theorem 1, indicating that distillation via $\mathcal{L}_{\text{rec-KD}}$ is sufficient.
Distillation weight.
@@
We sweep the distillation loss weight $\beta$ and observe best performance at $\beta=0.2$ (see Table 4). Too little weight underutilizes the teacher signal, while too much harms diversity by forcing alignment too strongly.
Conclusions
@@
In this work, we introduced SlotMatch, a simple and effective framework for distilling slot-based object representations from large teacher models into lightweight students. By aligning slots directly via cosine similarity, our method avoids auxiliary objectives. We showed both theoretically and empirically that this objective is sufficient to transfer semantic structure, leading to a student that outperforms its teacher in segmentation quality, while being significantly more efficient. Our results on MOVi-E and YTVIS established new state-of-the-art performance among unsupervised slot-based video models. In future work, we aim to continue our research in scaling down object-centric models for real-world deployment.
Supplementary
@@
In the supplementary, we include the demonstration for Theorem 1, additional quantitative and qualitative results, reproducibility details, and the reproducibility checklist.
Theoretical Demonstration
We provide the proof of Theorem 1 below.
@@
By replacing the definition in Eq. (6) in Eq. (4), we obtain the following:
@@
  • If the slots are normalized (i.e. $r=1$), the bound simplifies to:
@@
Additional Results
Results on additional dataset.
To further confirm that our framework generalizes to other datasets, we report additional results on the DAVIS 2017 dataset (Caelles et al. 2019). Our option is motivated by the fact that DAVIS contains high-quality annotations and challenging video sequences featuring complex object motion, occlusion, and dynamic backgrounds. DAVIS consists of short video clips with dense pixel-wise instance masks annotated per frame, making it particularly well-suited for evaluating fine-grained object-centric representations.
In Table 5, we compare SlotMatch with its SlotContrast teacher and a student that is trained from scratch (via SlotContrast without distillation), on the DAVIS dataset. The results further confirm the observations from the main article, namely that SlotMatch produces superior performance levels than the SlotContrast teacher, while also providing clear benefits over the student trained from scratch.
Method Image Video
FG-ARI $\uparrow$
mBO $\uparrow$
FG-ARI $\uparrow$
mBO $\uparrow$
SlotContrast (Teacher) 89.1 12.1 73.3 11.8
Student (no KD) 93.7 11.9 76.6 11.6
SlotMatch (ours) 95.6 14.9 92.8 14.7
Table 5: Comparison of video object segmentation performance on DAVIS 2017. We report FG-ARI and mBO scores at both image and video levels.
Ablation for Hungarian matching.
As slots are inherently permutation-invariant, there is no guarantee for one-to-one alignment between student and teacher slots. This presents a challenge for direct knowledge transfer. To this end, we explore two matching strategies: (i) aligning slots by index positions (as proposed in the main article) or (ii) using the Hungarian algorithm to find the most similar slot pairs based on their features, before applying the loss. As shown in Table  6, distillation without Hungarian matching yields better performance in our setup, while also being more computationally efficient. A likely cause for this conclusion is that strict matching over-constrains the learning process, negatively impacting the convergence of the distillation model and leading to suboptimal results. Therefore, all results reported in the main paper omit the Hungarian matching step.
Method Image Video
FG-ARI $\uparrow$
mBO $\uparrow$
FG-ARI $\uparrow$
mBO $\uparrow$
SlotContrast (Teacher) 83.9 32.4 81.7 28.6
Student (no KD) 80.1 28.9 76.6 30.1
SlotMatch (w/ matching) 83.5 31.8 81.5 28.6
SlotMatch (w/o matching) 84.1 33.6 81.8 30.5
Additional qualitative results.
In Figure 4, we present a qualitative comparison of segmentation results produced by the teacher, student model trained without distillation, and our SlotMatch model, across MOVi-E (left-hand side) (Ghorbani et al. 2021) (left) and YTVIS-2021 (Yang et al. 2021b) (right-hand side).
Hyperparameter MOVi-E YTVIS
Teacher Student Teacher Student
Backbone ViT-B/14 ViT-S/14 ViT-B/14 ViT-S/14
Feature size ($m$) 768 384 768 384
Slot dim ($S$) 128 128 64 64
#Slots ($N$) 15 15 7 7
Input size 3362 3362 5182 5182
#Patches 576 576 1369 1369
Batch size 16 16 64 64
Learning rate 0.0004 0.0004 0.0008 0.0008
Total steps 300K 300K 100K 100K
Loss weights (1.0, 0.5)
@@
Notably, green boxes indicate new or corrected slots introduced by our method, while red boxes highlight limitations of both the teacher and non-distilled student. These visual improvements confirm that distillation from the slots of a strong teacher not only preserves the original slot grouping, but also enhances the student’s robustness to occlusion, lighting variation, and clutter, all in an unsupervised fashion.
Reproducibility Details
@@
All experiments were conducted using two NVIDIA A100 GPUs (each with 40GB of VRAM). Training the student model on MOVi-E for 300K steps took approximately  138 hours. For YTVIS-2021, training required  39 hours for 100K steps. All models were implemented in PyTorch Lightning. Total GPU hours for each experiment are summarized in Table 8.
@@
Table 8: Approximate training time per experiment on two A100 GPUs.
@@
We used public benchmark datasets: MOVi-E (Ghorbani et al. 2021), YTVIS-2021 (Yang et al. 2021b), and OVIS (Qi et al. 2022). All datasets are publicly available and can be downloaded from the respective official repositories. Preprocessing for MOVi-E follows the original $\times336$ center crop and normalization to [1,1][-1,1]. For YTVIS, frames are resized using short-side resizing to 518 pixels with central cropping. OVIS is used for evaluation only. Temporal chunks of 4 frames are sampled per video for both training and evaluation. The corresponding preprocessing and data-related details are present in the configuration files.
@@
We will release pretrained weights for both teacher and student models on MOVi-E and YTVIS in our code repository [^1] https://github.com/dianagrigore/SlotMatch. Checkpoints will be provided with instructions for loading and evaluation. The code is released under the CC-BY-NC 4.0 license, and all dependencies are listed in the environment.yml file. The repository includes inference scripts, training pipelines, and evaluation tools for FG-ARI and mBO, along with an example configuration to reproduce results.

### v3

@@ 4.1 Datasets
MOVi-E. The Multi-Object Video (MOVi-E) dataset [ghorbani2021movi] is generated using the Kubric simulator, which provides ground-truth object segmentation for precise evaluation. MOVi-E includes up to 23 objects per scene and is filmed with linear camera motion. The dataset includes complex object interactions, occlusions, and realistic textures. The videos have a 24-frame length and a $\times256$ resolution. We use the official train and validation splits.
YTVIS-2021. To evaluate scalability to real-world scenarios, we use the YouTube Video Instance Segmentation 2021 (YTVIS-2021) dataset [vis2021]. It contains unconstrained real-world videos, capturing diverse scenes including indoor/outdoor environments, multiple object categories, and complex interactions. Videos vary in length (up to 76 frames) and resolution. We resize them to $\times518$ pixels. We report results on the official split.
DAVIS 2017. We report results on the DAVIS 2017 dataset [Perazzi-CVPR-2016] to evaluate fine-grained segmentation quality. This benchmark provides videos with dense temporal annotations, with every frame labeled using pixel-accurate object masks. The videos span diverse real-world scenarios, including people, animals, vehicles, and sports activities, capturing challenging phenomena like fast motion and cluttered backgrounds.
@@
4.2 Evaluation Metrics
We use two complementary evaluation metrics, namely the Foreground Adjusted Rand Index (FG-ARI) and the mean Best Overlap (mBO), and apply them at both image and video levels. FG-ARI measures how well the model groups pixels belonging to the same object. With a range of 0 to 1, and a higher value indicating better performance, its focus is on object discovery quality, i.e. if the model correctly identifies which pixels belong to the same object. FG-ARI compares predicted object masks against ground-truth segmentation masks. mBO measures the spatial precision of object masks, i.e. how accurately the predicted masks align with ground-truth boundaries. For each predicted mask, it finds the ground-truth mask with the highest IoU, then averages the IoU across all objects. mBO focuses on segmentation mask quality, measuring how precise and well-defined the predicted object boundaries are.
@@ 4.2 Evaluation Metrics
4.3 Baselines and Hyperparameters
Baselines. We compare SlotMatch against a range of recent object-centric video segmentation models. These include slot-based methods such as SAVi [kipf-ICLR-2022], STEVE [Singh-NeurIPS-2022], SOLV [Aydemir-NeurIPS-2023], VideoSAUR [Zadaianchuk-NeurIPS-2023], and its DINOv2-enhanced variant, VideoSAURv2 [Manasyan-CVPR-2025]. We reproduce the results of SlotContrast [Manasyan-CVPR-2025], which serves as the teacher in our framework based on knowledge distillation. We further introduce several new baselines, each employing a different distillation strategy, as follows:
  • DIOD [kara2024diod] employs self-distillation to continuously improve motion-guided object discovery via an iterative pseudo-labeling framework.
  • DINOSAUR [Seitzer-arxiv-2022] trains a student with a particular form of bottleneck that condenses the high-dimensional information from teacher features.
@@ 4.3 Baselines and Hyperparameters
Hyperparameters. We discuss reproducibility details and hyperparameter settings in the supplementary material.
4.4 Results
Quantitative results. In Table 1, we present comparative results on MOVi-E, YTVIS-2021 and DAVIS. Remarkably, SlotMatch consistently outperforms all prior methods across these datasets, including its teacher, SlotContrast [Manasyan-CVPR-2025]. Despite having $3.6\times$ fewer parameters and up to $2.7\times$ faster inference speed than SlotContrast (as per Table 2), our student achieves higher mask precision (mBO) and comparable or better object separation (FG-ARI). Notably, we improve video mBO on YTVIS from 32.4 (for SlotContrast) to 32.8 (for SlotMatch), while reducing latency by nearly half, showing that accurate object-centric segmentation does not require large models.
We find that mBO improvements are very large on the real-world DAVIS dataset, where the teacher may overfit to slot assignment, but underperform on precise masks. Interestingly, our student avoids this overfitting and better balances separation and alignment. This aligns with our hypothesis that slot supervision can offer better generalization than end-to-end training with larger capacity alone.
Qualitative results. In Figure 3, we present a qualitative comparison of segmentation results produced by the teacher, the student model trained without distillation, and our SlotMatch model, across MOVi-E [ghorbani2021movi] (left-hand side) and YTVIS-2021 [vis2021] (right-hand side).
Green boxes indicate new or corrected slots introduced by our method, while red boxes highlight limitations of both the teacher and non-distilled student. These visual improvements confirm that distillation from the slots of a strong teacher not only preserves the original slot grouping, but also enhances the student’s robustness to occlusion, lighting variation, and clutter, all in an unsupervised fashion.
Table 2: Comparison of inference speed, number of parameters, and GFLOPs between the SlotContrast teacher and our SlotMatch student. Measurements are performed on three GPUs (NVIDIA A100 with 40GB VRAM, NVIDIA 3090 with 24GB VRAM, NVIDIA 2080 with 12GB VRAM), using video samples from the YTVIS-2021 dataset.
Method #Params #GFLOPs $\downarrow$ Inference Time (ms) $\downarrow$
(M) $\downarrow$   A100   30903090   20802080
Teacher 91.0 8825.7 347.4 700.5 916.5
SlotMatch 25.2 3266.5 186.0 261.4 331.0
@@ 4.4 Results
[…]
Table 3: Zero-shot generalization results on the OVIS dataset, using models trained only on YTVIS (see performance on original dataset in Table 1). We compare the student trained from scratch (no KD) versus our SlotMatch student.
@@
Zero-shot generalization. In Table 3, we compare SlotMatch against the student trained from scratch (without distillation), in the zero-shot setup on the challenging OVIS dataset. Notably, slot-match maintains competitive video-level performance (FG-ARI: 34.8 vs. 34.6) compared with the student trained from scratch, while preserving slightly better spatial mask quality (mBO: 25.5 vs. 24.9). This demonstrates that distilling structured slot representations enhances robustness to occlusion and domain shift.
4.5 Ablation Studies
We conduct ablations to isolate the contribution of each component in our framework. The ablation results are summarized in Tables 1, 4, and 5.
Slot vs. feature and reconstruction KD. To test whether slot representations are the best target for distillation, we compare our method with variants that distill encoder or decoder features via MSE (see Table 1). While feature KD and reconstruction KD bring modest gains over the student (no KD) baseline, they remain below our slot-based formulation, by significant margins. This supports our hypothesis that distilling structured object-centric representations leads to better temporal consistency and segmentation quality.
Cosine vs. MSE loss. We compare the proposed loss based on cosine similarity with an alternative loss based on MSE over slot representations (see Table 1). Cosine-based distillation yields more stable improvements across datasets, especially on YTVIS, likely due to its scale invariance and better alignment with semantic structure in high-dimensional spaces.
@@ 4.5 Ablation Studies
Method Image Video
FG-ARI $\uparrow\,$
mBO $\uparrow$
FG-ARI $\uparrow$
mBO $\uparrow$
SlotContrast (Teacher) 83.9 32.4 81.7 28.6
Student (no KD) 80.1 28.9 76.6 30.1
SlotMatch (w/ matching) 83.5 31.8 81.5 28.6
SlotMatch (w/o matching) 84.1 33.6 81.8 30.5
@@
With or without reconstruction KD. To empirically determine if the proposed $\mathcal{L}_{\text{slot-KD}}$ loss is sufficient or not, we carry out experiments with an enhanced version of SlotMatch, where the distillation is also applied over reconstructed features via the $\mathcal{L}_{\text{rec-KD}}$ loss (see Table 1). Perhaps surprisingly, this double distillation procedure degrades performance by considerable margins on MOVi-E. The results show that adding additional distillation losses can make the distillation more difficult to tune, eventually leading to inferior results. In summary, the empirical results confirm Theorem 1, indicating that distillation via $\mathcal{L}_{\text{slot-KD}}$ is sufficient.
@@
Distillation weight. We sweep the distillation loss weight $\beta$ and obtain the best performance at $\beta=0.2$ (see Table 4). Too little weight underutilizes the teacher signal, while too much harms diversity by forcing alignment too strongly.
Hungarian vs. direct index matching. As slots are inherently permutation-invariant, there is no guarantee for one-to-one alignment between student and teacher slots. This presents a challenge for direct knowledge transfer. To this end, we explore two matching strategies: (i) aligning slots by index positions (proposed option) or (ii) using the Hungarian algorithm to find the most similar slot pairs based on their features, before applying the loss. As shown in Table 5, distillation without Hungarian matching yields better performance in our setup, while also being more computationally efficient. A likely cause for this outcome is that strict matching over-constrains the learning process, negatively impacting the convergence of the distillation model and leading to suboptimal results. Therefore, all results reported in the experiments omit the Hungarian matching step.
5 Conclusion
@@
In this work, we introduced SlotMatch, a simple and effective framework for distilling slot-based object representations from a large teacher model into a lightweight student. By aligning slots directly via cosine similarity, our method avoids auxiliary objectives. We showed both theoretically and empirically that this objective is sufficient to transfer semantic structure, leading to a student that outperforms its teacher in segmentation quality, while being significantly more efficient. Our results on MOVi-E, YTVIS and DAVIS established new state-of-the-art performance levels among unsupervised slot-based video models. In future work, we aim to continue our research in scaling down object-centric models for real-world deployment.
@@
In the supplementary, we include the demonstration for Theorem 1, additional quantitative and qualitative results, as well as reproducibility details.
Appendix A Theoretical Demonstration
We next provide the proof of our theorem.
@@ Appendix A Theoretical Demonstration
By replacing the definition from Eq. (6) in Eq. (4), we obtain the following:
@@
Table 6: Video segmentation results on DAVIS [Perazzi-CVPR-2016] with two different teacher models, VideoSAUR [Zadaianchuk-NeurIPS-2023] and VideoSAURv2 [Manasyan-CVPR-2025]. The top two scores for each metric are highlighted in blue bold (top method), orange bold (second best).
Method Image mBO $\uparrow$ Video mBO $\uparrow$
VideoSAUR [Zadaianchuk-NeurIPS-2023] (teacher) 23.9 18.4
Student (no KD) 17.4 14.4
SlotMatch (ours) 30.3 23.2
VideoSAURv2 [Manasyan-CVPR-2025] (teacher) 23.0 19.2
Student (no KD) 22.9 18.2
SlotMatch (ours) 31.2 28.9
@@
  • If the slots are normalized (i.er=1r=1), the bound simplifies to:
@@
Appendix B Additional Results
Qualitative results. In Figure 4, we showcase qualitative comparisons of segmentation masks on challenging examples from MOVi-E and YTVIS. Our method produces sharper and more temporally-consistent masks than the student model without distillation. On MOVi-E, SlotMatch segments overlapping or partially occluded objects more cleanly, while on real-world YTVIS data, it shows improved boundary alignment and fewer slot collisions. These results highlight the benefit of slot-level supervision in guiding the student towards focusing on meaningful object structure.
Table 7: Image segmentation results on MS COCO [Lin-ECCV-2014] with DINOSAURv2 [Seitzer-arxiv-2022] as teacher. The top two scores for each metric are highlighted in blue bold (top method), orange bold (second best).
Method Image FG-ARI $\uparrow$ Image mBO $\uparrow$
DINOSAURv2 [Seitzer-arxiv-2022] (teacher) 44.7 29.3
Student (no KD) 41.5 28.6
SlotMatch (ours) 42.3 28.8
Results with different teachers. In Table 6, we report results with additional teacher models, namely VideoSAUR [Zadaianchuk-NeurIPS-2023] and VideoSAURv2 [Manasyan-CVPR-2025], on the DAVIS [Perazzi-CVPR-2016] dataset. The SlotMatch student based on DINOv2 is compared with an equivalent architecture without distillation (no KD), as well as its corresponding teacher model. The results follow the same trend as in the main experiments, i.eSlotMatch consistently outperforms both the student (no KD) and teacher models. These results confirm that SlotMatch generalizes across multiple teacher models.
Hyperparameter MOVi-E YTVIS-2021 DAVIS 20217
Teacher Student Teacher Student Teacher Student
Backbone ViT-B/14 ViT-S/14 ViT-B/14 ViT-S/14 ViT-B/14 ViT-S/14
Feature size ($m$) 768 384 768 384 768 384
Slot dim ($S$) 128 128 64 64 128 128
#Slots ($N$) 15 15 7 7 4 4
Input size $\times336$ $\times336$ $\times518$ $\times518$ $\times336$ $\times336$
#Patches 576 576 1369 1369 576 576
Batch size ($B$) 16 16 64 64 16 16
Learning rate ($\eta$) 0.0004 0.0004 0.0008 0.0008 0.0008 0.0008
Total steps 300K 300K 100K 100K 100K 100K
Loss weights ($\alpha,\beta$) (1, 0.2)
@@ Appendix B Additional Results
Image segmentation results. In Table 7, we present image segmentation results on the MS COCO [Lin-ECCV-2014] dataset, employing DINOSAURv2 [Seitzer-arxiv-2022] as teacher. SlotMatch surpasses the analogous model (without distillation), but remains below the teacher. We conclude that SlotMatch can obtain competitive results in image segmentation, with a reduced computational cost. Yet, the true benefit of SlotMatch is more obvious in video segmentation, where its ability to maintain temporal consistency of slots plays a key role.
Appendix C Reproducibility Details
@@
All experiments were conducted using two NVIDIA A100 GPUs (each with 40GB of VRAM). Training the student model on MOVi-E for 300K steps took approximately 138 hours. For YTVIS-2021, training required about 39 hours for 100K steps. All models were implemented in PyTorch Lightning. Total GPU hours for each experiment are summarized in Table 9.
Table 9: Approximate training time per experiment on two A100 GPUs.
@@
DAVIS 100K $\approx$ 25
@@
We used public benchmark datasets: MOVi-E [ghorbani2021movi], YTVIS-2021 [vis2021], DAVIS [Perazzi-CVPR-2016] and OVIS [qi2022occluded]. All datasets are publicly available and can be downloaded from the respective official repositories. Preprocessing for MOVi-E follows the original $\times336$ center crop and normalization to [1,1][-1,1]. For YTVIS, frames are resized using short-side resizing to 518 pixels with central cropping. OVIS is used for evaluation only. Temporal chunks of 4 frames are sampled per video for both training and evaluation. The corresponding preprocessing and data-related details are present in the configuration files.
@@
We will release pretrained weights for both teacher and student models on all datasets in our code repository. Checkpoints will be provided with instructions for loading and evaluation. The code is released under the CC-BY-NC 4.0 license, and all dependencies are listed in the environment.yml file. The repository includes inference scripts, training pipelines, and evaluation tools for FG-ARI and mBO, along with an example configuration to reproduce results.